OpenStudy (anonymous):

Find an equation of the line passing through the given points. Use function notation to write the equation. (4,-22) and (-8,42)

hero (hero):

Use y = mx + b 42 = -8m + b -22 = 4m + b b = 42 + 8m b = -22 - 4m b = b 8m + 42 = -22 - 4m 8m + 4m = -22 - 42 12m = -64 m = -64/12 m = -16/3 b = 42 + 8m b = 42 + 8(-16/3) b = 42 -128/3 b = 136/3 - 128/3 b = (136 - 128)/3 b = 8/3 y = -16/3(x) + 8/3 Change y to f(x): f(x) = -16/3(x) + 8/3

OpenStudy (anonymous):

Thanks